# Reconstruction-based Multi-Normal Prototypes Learning for Weakly Supervised Anomaly Detection

### v1

@@ 5 Conclusions
In conclusion, our proposed anomaly detection method demonstrates significant improvements over existing techniques by effectively leveraging a small amount of labeled anomaly data alongside a large amount of unlabeled data. Our method’s robustness to anomaly contamination and its ability to generalize to unseen anomalies make it highly suitable for real-world applications across various domains. Future work could explore extending our method to other types of data, such as images, to further validate its versatility and effectiveness. For the reconstruction learning component, we could also consider using more advanced variants like Variational AutoEncoders (VAEs) to enhance feature extraction capabilities.

### v2

@@ V Conclusions
In conclusion, our proposed anomaly detection method demonstrates significant improvements over existing techniques by effectively leveraging a small amount of labeled anomaly data alongside a large amount of unlabeled data. Our method’s robustness to anomaly contamination and its ability to generalize to unseen anomalies make it highly suitable for real-world applications across various domains, including network security[6], financial fraud detection[7], and medical diagnostics[9]. Future work could explore extending our method to other types of data, such as images, to further validate its versatility and effectiveness. For the reconstruction learning component, we could also consider using more advanced variants like Variational AutoEncoders (VAEs)[49] to enhance feature extraction capabilities.